INAF LBC+MODS queue observing and AO PISCES, 2012 June 16-17

Observer: C. Arcidiacono, R. Briguglio, K. Boutsia, M. Faccini, F. Vitali
Telescope Operator: S. Allanson
Support Astronomer: D. Thompson
Instrument: LBC + PISCES

Summary

Thick clouds and rain during the first part of the night. The dome remains closed.
At around 6:40UT the sky clears and we open. It is too late to do the LBC target and we move on to Pisces.
There were passing clouds during the whole night, so we had to do bright targets. The seeing was variable: ~1.5" at the beginning of the night going down to 0.9" towards the end.
The AO system did not give us significant problems.
The main problem we had with the observe routines, was that it occasionally skips images with the extended header (it does not save them in /newdata). Doug Miller was informed and will work on the problem during daytime.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Barbieri KIC4060 H,J,Ks ~10min X filter + Sky COMPLETED
Barbieri KIC6448 H,J,Ks ~10min X filter + Sky COMPLETED

Details

06:40 Opening Dome
  • too late to do LBC targets, configuring for pisces

07:00 Slewing at the Cesaroni target.
  • dimm seeing 1.6"
  • they are not able to close the loop on the pointing star
  • trying to close the loop on the reference star

07:08 forgot to turn off the treehouse fans - abort preset
  • telescope to the zenith and C.A. goes to the dome

07:19 resend preset - Cesaroni
  • loop closed but we have at the center the reference star (test)
  • 07:28 - the loop opened ... probably because the seeing is not that good...

07:30 resend preset with offset (before the ref. star was at the center) in order to start taking data...
  • when we send the filter command - it changes the offset in z and the loop opens
  • it happened again just like before!!

07:41 resend preset
  • dimm seeing 1.3"
  • loop open again - there are probably clouds passing through because this time the open loop was caused by a big change in flux.

07:55 AdSec RIP

08:00 resend preset - Cesaroni
  • no way to close the loop on this target because of the clouds

08:10 we go to a brighter target - Barbieri
  • send preset KIC4060
  • we notice that not all images are saved in newdata with the long header - call DougMiller
  • in the meantime we keep obtaining images

09:07 the loop opened because of a failed "resume" during the sky offset
  • resend preset

09:23 loop open after a "skip frame" - clouds??
  • we ask the T.O. to offset the telescope to get the sky frames

09:28 slewing to Cesaroni to try again...
  • it is not visible because of clouds

09:35 slewing to KIC6448

10:50 need to rebbot the bcu

11:08 slew to Moscadelli
  • the reference star is at the edge of the aquire ccd - it does not manage to run acq_ref
  • we just send the preset and will try to acquire the ref star manually
  • did not manage to acquire the star - too faint (we are well within twilight)
  • actually there was a filter (silver) in the ccd47 ... when they put clear, they manage to see the star

11:25 we slew for flats
  • mostly Ks because we have H and J from yesterday

12:30 send darks

-- KonstantinaBoutsia - 17 Jun 2012
Topic revision: r4 - 22 Jun 2012, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback